1. The utility function for an individual is given by the equation: π‘ˆ = 𝑋1 0.25𝑋2 0.75 and their budget constraint is 𝑃1𝑋1 + 𝑃2𝑋2 = 𝑀 (i) Using the Lagrange, and showing all work, Derive the demand functions for good 𝑋1 π‘Žπ‘›π‘‘ π‘”π‘œπ‘œπ‘‘ 𝑋2 (ii) Given that the price for good 1, p1= k2.5 and the price for good 2, p2=k5, and the consumer’s income is k100. Calculate and graphically present the optimal choice for good 1 and good 2


1
Expert's answer
2021-10-20T09:58:01-0400

i

The Lagrangian is set up as follows :-

"L = X_1^{0.25} X_2^{0.75} + \u03bb(M - P_1X_1 + P_2X_2 )"


First order conditions :-

"\frac{\delta L}{\delta X_1}=0.25(\frac{X_2}{X_1})^{0.75}-\lambda P_1=0.................(1)"


"\frac{\delta L}{\delta X_2}=0.75(\frac{X_1}{X_2})^{0.25}-\lambda P_2=0.................(2)"


"\frac{\delta L}{\delta \lambda}=M-P_1X_1-P_2X_2=0.................(3)"


"\frac{(1)}{(2)}" gives


"\frac{0.25(\frac{X_2}{X_1})^{0.75}}{0.75(\frac{X_1}{X_2})^{0.25}}=\frac{P_1}{P_2}"


simplifying


"X_2=\frac{3P_1X_1}{P_2}"


Now, substituting the value of X2Β in (3) gives -


"M-P_1X_1-P_2(\frac{3P_1X_1}{P_2})=0"


solving


"X_1=\frac{M}{4P_1}"


Substituting this value of X1Β in the previously found value of X2, we get -


"X_2=\frac{3M}{4P_2}"


These are the required demand functions.


ii

(ii) Given:

P1Β = k 2.5

P2Β = k 5

M = k 100

So XΒ 1 and X2 Β becomes -


"X_1=\frac{100}{4\times2.5}=10"


"X_2=\frac{3\times 100}{4\times 5}=15"


The budget constraint becomes -

2.5X1Β + 5X2Β = 100


Graphical representation :-





Here, the green line shows the budget line while the pink curve shows the indifference curve. The intercepts of the budget line have been obtained as -


"X1 = \frac{100}{2.5} = 40
\\
X2 = \frac{100}{5} = 20"


E is the optimal bundle of the individual.